Soybean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280-lb mixture that is 10% protein?
Answer by mananth(16949)

Soybean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280-lb mixture that is 10% protein?
...
Soybean meal is 14% protein x lb
cornmeal is 7% protein. 280-x lb
10% protein 280-lb mixture
...
14x+7(280-x)=10*280
14x+1960-7x=2800
7x+1960=2800
-1960
7x=840
/7
x=120 lb. ----------soyabean
160 lb -------------cornmeal
...
m.ananth@hotmail.ca